# Simulating and investigating various dynamic aspects of $\rm{H}_{2}\rm{O}$-related hydrogen bond model

### v2

@@ Abstract
A simple $\rm{H}_{2}\rm{O}$-related hydrogen bond model, modified from the Jaynes–Cummings model, is proposed and its various dynamic aspects are investigated theoretically. In this model, the formation and breaking processes of hydrogen bond are accompanied by the creation and annihilation of the thermal phonon of the medium. A number of simplifying assumptions about the dynamics of the molecules involved are used. Rotating wave approximation is applied under consideration of the strong-coupling condition. Dissipative dynamics under the Markovian approximation is obtained through solving the quantum master equation — Lindbladian. The probabilities of reaction channels involving hydrogen bond depending on the parameters of the external environment, are obtained. Differences between unitary and dissipative evolutions are discussed. Consideration is given to the effect of all kinds of potential interactions and dissipations on evolution. Consideration is also given to the reverse processes (inflows) of dissipations. The results show that the magnitude changes of the interactions and dissipations have slight effect on the formation of hydrogen bond, but the variation of the inflows significantly affects the formation of hydrogen bond. According to the findings, the dynamics of $\rm{H}_{2}\rm{O}$-related hydrogen bond model can be controlled by selectively choosing system parameters. The results will be used as a basis to extend the research to more complex chemical and biological model in the future.
hydrogen bond, finite-dimensional quantum electrodynamics, Markovian open system, water molecule, thermal phonon
@@ I Introduction
Complex quantum system modeling is one of the most important directions in computational mathematics today, especially in the computational fields involving polymer chemistry and macromolecular biology [1, 2, 3]. A computer chemistry simulator with predictive power requires an entirely quantum treatment; its creation poses a serious challenge to computational mathematics due to the exponentially growing complexity of calculations — curse of dimensionality [4, 5]. The dynamics of chemical transformations, in contrast to the molecular dynamics of ready-made molecules, requires the involvement of an electromagnetic field, which further aggravates the problem of complexity. The most important type of chemical transformations is the formation and disintegration of hydrogen bonds between molecules, which are responsible for the formation and disintegration of macromolecules. Such bonds are formed by a proton tunneling between two conventional potential wells between two molecules. The discovery of hydrogen bonds is attributed to T.S. Moore and T.F. Winmill [6], and the description of hydrogen bonding in water was first described in 1920 [7]. Hydrogen bonds are much weaker than covalent bonds (in a water ($\rm{H}_{2}\rm{O}$)2 dimer, the energy of a hydrogen bond is only an order of magnitude higher than the thermal energy at room temperature, while for a covalent bond in an $\rm{OH}$ molecule the energy is 200 times greater) and therefore their formation and decay are easily controlled by external influences, for example, temperature serves as a mechanism for the transformation of macromolecules. Such transformations occur, for example, during the synthesis of DNA, the double helix of which is connected precisely by these bonds. Hydrogen bonds in water are responsible for its extreme heat capacity; their short lifetime — about $10^{-11}$ seconds [8], determines the flexibility of water clusters and their good interaction with donor molecules [9]. In recent years, hydrogen bonds have become one of the main objects of research into quantum processes related to biology. Decoherence in hydrogen bonds was considered in [10]. The entangled spin states that arise in them are in [11]. A more chemical consideration of the hydrogen bonds that arise in the $\alpha$-helix of proteins participating in the protein machinery of living organisms is presented in [12]. Hybrid bonds in liquids, including proton tunneling, as well as in water clusters, were studied in [13, 14]. The possibility of using proton tunneling to recognize molecules was also explored in [15]. The chemical role of hydrogen bonding in enzymes was studied in [16].
@@
Consideration of this type of bond involves many elements. In our work, a highly simplified model of hydrogen bond that can be easily scaled to complex molecular systems to make their simulation possible on modern computers, is proposed. A key contribution of this paper is the cavity quantum electrodynamics (QED) models [17, 18, 19, 20, 21], which are easy to implement in the laboratory and offers a unique scientific paradigm for studying light–matter interaction. The cavity QED model includes the Jaynes–Cummings model (JCM) [22] and the Tavis–Cummings model (TCM) [23] as well as their generalizations [24]. Many studies have been conducted recently in the field of these models, including those on quantum gates [25, 26], quantum many-body phenomena [27], entropy [28], quantum discord [29], dark states [30, 31, 32, 33, 34, 35, 36, 37], phase transitions [38, 39], etc [40, 41, 42, 43, 44, 45, 46, 47]. As a basis, the generally accepted JCM is introduced and modified appropriately so that the presence of a hydrogen bond will play the role of the ground state of (conditional) atoms, and its absence will play the role of the excited states of the atoms. The optical cavity will correspond to the region where the emitted phonon can be again absorbed by the molecular structure with the destruction of the resulting hydrogen bond.
@@
This paper is organized as follows. The $\rm{H}_{2}\rm{O}$-related hydrogen bond model is proposed in Sec. II. After introducing the physico-biological mechanisms of the target model in Sec. II.1, its Hilbert space and Hamiltonian is constructed in Sec. II.2 and quantum master equation (QME) is introduced in Sec. II.3. The numerical method is introduced in Sec. III. The results of our numerical simulations is presented in Sec. IV, including comparison between unitary and dissipative evolutions in Sec. IV.1, and the effects of interactions, dissipations and reverse processes (inflows) of dissipations on the evolution in Secs. IV.2$\sim$ IV.4. Besides, the effect of external impulses on evolution is shown in Sec. IV.5. Some brief comments on our results in Sec. V close out the paper. Some technical details are included in Appendices A and B.
@@ II Hydrogen bond model
The hydrogen bond dynamics in media as a dynamics of polariton of the group of real particles — two molecules and quasiparticles (photons and phonons), is represented. Strictly speaking, the quantum state of polariton has the form
$\lambda_{photon}|photon\rangle+\lambda_{phonon}|phonon\rangle+\lambda_{particle}|particle\rangle$ (1)
This is hardly possible to deal with this form, because any interaction (photon–phonon, phonon–atom, photon–atom) represents nontrivial task. Thus, the initial photon can transform to a few phonons. So it is accepted to call the polariton of phonons if there is no explicit photons in advance. This terminology is followed and the system in the framework of Jaynes–Cummings scheme is presented with Hamiltonian
$a^{{\dagger}}\sigma+a\sigma^{{\dagger}}+a^{{\dagger}}a+\sigma^{{\dagger}}\sigma$ (2)
@@
The weakness of the hydrogen bond causes its strong dependence on external conditions, in particular, on the ambient temperature. The temperature itself can be conventionally represented as a graph of the average number of thermal phonons versus their frequency. In [48], the temperature effect on atoms within the Jaynes–Cummings model was represented as terms of the Hamiltonian of the form
$g(a_{\omega}^{\dagger}+a_{\omega})\sigma_{\omega}^{\dagger}\sigma_{\omega}$ (3)
where $a_{\omega}$ is the annihilation operator of the phonon of the selected mode $\omega$, $\sigma_{\omega}$ is the relaxation operator of the atom. This approach is extremely computationally expensive, since it requires explicit inclusion of phonons in the basic states of the model, which immediately causes a huge increase in the required memory and does not allow scaling the model to multi-molecular structures to study new collective effects. In addition, the case of atomic excitations is very different from Eq. (3), because their energy is several orders of magnitude greater than the energy of a single phonon, so that a term of the form Eq. (3) corresponds only to dephasing, but not to direct interaction of phonons with matter.
The decoherence is described caused by the influence of the medium using the QME with a decoherence factor in the form of phonon annihilation. This approach is simpler and more efficient than an independent a-priori introduction of decoherence with a Gaussian factor in [10], and also allows for simple scaling.
Interest is only given to the frequency that most strongly affects the hydrogen bond. This influence can be represented as a term in the Hamiltonian of the form
$H_{int}=g_{hb}(a_{phn}^{{\dagger}}\sigma_{hb}+a_{phn}\sigma_{hb}^{\dagger})$ (4)
where $\sigma_{hb}$ is the operator of hydrogen bond formation accompanied by phonon emission, and $a_{phn}$ is the operator of phonon absorption leading to bond decay. Direct inclusion of such a term in the Hamiltonian is also unacceptable for us, since explicit phonons again arise in the basic states. An approximation in the form of the average number of phonons $n_{av}$ at the resonance frequency sensitive to hydrogen bonding is applied; since the operators of phonon annihilation and creation are proportional to the square root of their number $\sqrt{n_{av}}$, $H_{int}$ is replaced with the operator
$H^{\prime}_{int}=g_{0}\sqrt{n_{av}}(\sigma_{hb}^{\dagger}+\sigma_{hb})$ (5)
Then the basic state will not contain an explicit number of phonons and the model is able to be scaled to complex molecular systems.
@@
${\cal G}\left(T\right)=c\sum\limits_{n=0}^{\infty}e^{\frac{-\hbar\omega n}{KT}}|n\rangle\langle n|$ (6)
where $\hbar$ is reduced Planck constant, $\omega$ is phononic mode, $n$ is phonon number, $T$ is the cavity temperature at a given frequency mode $\omega$, $K$ is the Boltzmann constant, $c$ is normalization factor. $\gamma_{in}$ and $\gamma_{out}$, which are the intensities of the inflow and dissipation of phonons into and out of a notional cavity around the molecules, are proposed to form a geometric progression with the denominator [49]
$\mu=\frac{\gamma^{in}}{\gamma^{out}}=e^{\frac{-\hbar\omega}{KT}}$ (7)
where $\gamma^{in}$ refers to the overall spontaneous inflow rate, $\gamma^{out}$ refers to the overall spontaneous emission rate. A stable temperature occurs only when $\gamma_{in}<\gamma_{out}$. Knowing these coefficients, or knowing the temperature $T$ directly, $n_{av}$ is obtained. In practice, the coefficients are found only by optimizing them using neural networks based on the experimental results. The temperature at a fixed photonic mode is determined in a similar way. Photonic modes relate to transformations of electron states, and phononic modes relate to proton oscillations.
@@ II.1 The target model
Figure 1: (online color) Hydrogen bond formation mechanism. Scale model of two $\rm{H}_{2}\rm{O}$ molecules is shown in panel (a). Mechanism of ion attraction between hydrogen and oxygen atoms is shown in the form of a ball–and–stick model in panel (b). Brown circles (balls) represent oxygen atom, red circles (balls) represent hydrogen atom in excited state, black solid lines stand for covalent bond and gray dotted lines represent ion attraction; especially, black dashed line indicate stretched covalent bond and green ball in panel (a) represent a proton in ground state.
@@
Interaction between two water molecules causes micro-oscillations of the hydrogen atom in one of water molecules, it allows hydrogen bond to form between water molecules, but this process does not break covalent bond. The hydrogen bond formation mechanism is as shown in Fig. 1. In panel (a), when two $\rm{H}_{2}\rm{O}$ molecules that are moving freely are far apart, hydrogen bond cannot form between the molecules. However, when the molecules move closer together, the hydrogen atom (proton) of one molecule and the oxygen atom of another molecule attract each other and stable hydrogen bond is obtained. In panel (b), the hydrogen atom of one molecule “donor” is attracted by the oxygen atom of another molecule “receptor” to produce the tunneling effect. The proton tunneling in the normal state does not destroy the covalent bond between proton and its parent molecule, but only deforms it. In the case without considering of electrons, the effect of some stretching of the covalent bond on the hydrogen bond formation is ignored, and it is considered as a normal state with a covalent bond to avoid cluttering the target model. In the model with electrons, this difference is no longer ignored. Theoretically, in addition to 2 covalent bonds, for the oxygen atom, there are actually 2 more hydrogen bonds with protons from neighboring molecules. The oxygen atom and its covalent and hydrogen bonds form a tetrahedron.
@@
A hydrogen bond between water molecules occurs when they approach each other at a distance $|0\rangle_{dist}$, allowing one of the protons, covalently bonded to the parent molecule, to tunnel between it and another neighbouring molecule, while maintaining the existing covalent bond. The qubit $|d\rangle_{dist}$ describing the relative position of proton in the system, is introduced, and $d\in\{-1,0,1,2\}$. $dist=2$ means that the molecules are separated by a large distance, at which the formation of a hydrogen bond is impossible in principle. The state $dist=1$ means that the molecules are approaching each other at a critical distance, and although it does not allow the creation of a hydrogen bond, it allows the molecules to tunnel to the state $dist=0$, where the bond becomes possible. The state $dist=-1$ means the presence of a stable hydrogen bond, here it is an irreversible process. However, the influence of temperature on formation of hydrogen bond is investigated, because change of temperature will cause inflow of phonon. Thence, in this situation this process becomes reversible, that means the breaking of hydrogen bonds becomes possible.
@@
$H_{i}=H_{i}^{basis}+H_{i}^{cogerence}$ (8)
where $H_{i}^{basis}$ is diagonal in the standard basis, and $H_{i}^{cogerence}$ is not diagonal in this basis. The term $H_{i}^{basis}$ is added to the total Hamiltonian in any case, and $H_{i}^{cogerence}$ is added only if the condition $A$ is satisfied. This remark about the conditional Hamiltonian will be valid for further models as well. This method achieves equality of rest energies for all terms of the general Hamiltonian. So the condition operator $Cond$ means only the inclusion of coherent terms in the Hamiltonian. For example, the proton (and, subsequently, the electron) jumps only if $d=0$, that is, when the molecules are close. Only at a large distance between the molecules is it as if “frozen”, and does not lead to a change in the state of the proton (and, subsequently, the electron). And as soon as the molecules are close, the motions of these particles begin. The standard basis thus turns out to be distinguished from all others. In this work, a single hydrogen bond between a pair of water molecules, which will be extended to more hydrogen bonds using Eq. (8) in the future, is considered.
Hydrogen bonding is a fundamental interaction in chemistry and biology, playing a crucial role in the structure and function of a wide range of molecules, from water and small organic compounds to large biomolecules like DNA and proteins. Although there is no exact universal value for the time of hydrogen bond formation, it can be assumed that it can be in the range from femtoseconds to picoseconds, that is, from $10^{-15}$ to $10^{-12}$ seconds. This is due to the fact that the formation of hydrogen bonds involves the reorganization of electron clouds — a very fast quantum mechanical process. Time scale measurements of hydrogen bond dynamics can be performed using ultrafast spectroscopy techniques [50, 51, 52].

### v3

# Simulating and investigating various dynamic aspects of the $\rm{H}_{2}\rm{O}$-related hydrogen bond model

@@ Abstract
A basic model of hydrogen bonds related to $\rm{H}_{2}\rm{O}$, which is adapted from the Jaynes–Cummings model, is suggested, and its different dynamic features are studied theoretically. In this model, the making and breaking of hydrogen bonds happen alongside the creation and destruction of phonons in the surrounding medium. A number of simplifying assumptions about the dynamics of the molecules involved are used. The rotating wave approximation is applied under consideration of the strong-coupling condition. Dissipative dynamics under the Markovian approximation is obtained through solving the quantum master equation — Lindbladian. We obtain the probabilities of reaction channels involving hydrogen bonds based on the parameters of the external environment. Differences between unitary and dissipative evolutions are discussed. Consideration is given to the effects of all kinds of potential interactions and dissipation on evolution. Consideration is also given to the reverse processes (inflows) of dissipation. The results show that the magnitude changes of the interactions and dissipation have a slight effect on the formation of hydrogen bonds, but the variation of the inflows significantly affects the formation of hydrogen bonds. According to the findings, the dynamics of the $\rm{H}_{2}\rm{O}$-related hydrogen bond model can be controlled by selectively choosing system parameters. The results will be used as a basis to extend the research to more complex chemical and biological models in the future.
hydrogen bond, finite-dimensional QED, Markovian open system, water molecule, phonon
@@ I Introduction
Complex quantum system modeling is one of the most important directions in computational mathematics today, especially in the computational fields involving polymer chemistry and macromolecular biology [1, 2, 3]. A computer chemistry simulator with predictive power requires an entirely quantum treatment; its creation poses a serious challenge to computational mathematics due to the exponentially growing complexity of calculations — curse of dimensionality [4, 5]. The dynamics of chemical transformations, in contrast to the molecular dynamics of ready-made molecules, require the involvement of an electromagnetic field, which further aggravates the problem of complexity. The most important type of chemical transformation is the formation and disintegration of hydrogen bonds between molecules, which are responsible for the formation and disintegration of macromolecules. Such bonds are formed by a proton tunneling between two conventional potential wells between two molecules. The discovery of hydrogen bonds is attributed to T.S. Moore and T.F. Winmill [6], and the description of hydrogen bonding in water was first described in 1920 [7]. Hydrogen bonds are much weaker than covalent bonds (in a water ($\rm{H}_{2}\rm{O}$)2 dimer, the energy of a hydrogen bond is only an order of magnitude higher than the thermal energy at room temperature, while for a covalent bond in an $\rm{OH}$ molecule the energy is 200 times greater), and therefore their formation and decay are easily controlled by external influences; for example, temperature serves as a mechanism for the transformation of macromolecules. Such transformations occur, for example, during the synthesis of DNA, the double helix of which is connected precisely by these bonds. Hydrogen bonds in water are responsible for its extreme heat capacity; their short lifetime is about $10^{-11}$ seconds [8] — determines the flexibility of water clusters and their good interaction with donor molecules [9]. In recent years, hydrogen bonds have become one of the main objects of research into quantum processes related to biology. Decoherence in hydrogen bonds was considered in [10]. The entangled spin states that arise in them are in [11]. A more chemical consideration of the hydrogen bonds that arise in the $\alpha$-helix of proteins participating in the protein machinery of living organisms is presented in [12]. Hybrid bonds in liquids, including proton tunneling, as well as in water clusters, were studied in [13, 14]. The possibility of using proton tunneling to recognize molecules was also explored in [15]. The chemical role of hydrogen bonding in enzymes was studied in [16].
@@
Consideration of this type of bond involves many elements. In our work, we propose a highly simplified model of hydrogen bonds that can be easily scaled to complex molecular systems, making their simulation possible on modern computers. A key contribution of this paper is the cavity quantum electrodynamics (QED) models [17, 18, 19, 20, 21], which are easy to implement in the laboratory and offer a unique scientific paradigm for studying light–matter interaction. The cavity QED model includes the Jaynes–Cummings model (JCM) [22] and the Tavis–Cummings model (TCM) [23], as well as their generalizations [24]. Many studies have been conducted recently in the field of these models, including those on quantum gates [25, 26], quantum many-body phenomena [27], entropy [28], quantum discord [29], dark states [30, 31, 32, 33, 34, 35, 36, 37], phase transitions [38, 39], etc. [40, 41, 42, 43, 44, 45, 46, 47, 48, 49]. As a basis, the generally accepted JCM is introduced and modified appropriately so that the presence of a hydrogen bond will play the role of the ground state of (conditional) atoms, and its absence will play the role of the excited states of the atoms. The optical cavity will correspond to the region where the emitted phonon can be again absorbed by the molecular structure through the destruction of the resulting hydrogen bond.
@@
This paper is organized as follows. The $\rm{H}_{2}\rm{O}$-related hydrogen bond model is proposed in Sec. II. After introducing the physico-biological mechanisms of the target model in Sec. II.1, its Hilbert space and Hamiltonian are constructed in Sec. II.2, and the quantum master equation (QME) is introduced in Sec. II.3. The numerical method is introduced in Sec. III. The results of our numerical simulations are presented in Sec. IV, including a comparison between unitary and dissipative evolutions in Sec. IV.1, and the effects of interactions, dissipation, and reverse processes (inflows) of dissipation on the evolution in Secs. IV.2 $\sim$ IV.4. Besides, the effect of external impulses on evolution is shown in Sec. IV.5. Some brief comments on our results in Sec. V close out the paper. Some technical details are included in Appendices A and B.
@@ II Hydrogen bond model
The hydrogen bond dynamics in the medium are represented as a dynamics of polariton, which consists of a group of real particles (two molecules and photons) and quasi-particles (phonons). The reason why we introduce phonons is that the formation and breaking of hydrogen bonds belong to intermolecular interactions. This intermolecular vibration exists in the form of phonons in solid ice crystals; while for liquid water, although it does not have the long-range order characteristics of ice crystals, it still has short-range order characteristics within 1 to 5 atomic distances. At this time, the concept of local phonons or quasi-phonons can still be used to represent intermolecular vibrations. Strictly speaking, the quantum state of polariton has the following form
$\lambda_{photon}|photon\rangle+\lambda_{phonon}|phonon\rangle+\lambda_{particle}|particle\rangle.$ (1)
This is hardly possible to deal with in this form because any interaction (photon–phonon, phonon–atom, photon–atom) represents a nontrivial task. Thus, the initial photon can transform into a few phonons. So it is accepted to call the polariton of phonons if there are no explicit photons in advance. This terminology is followed, and the system in the framework of the Jaynes–Cummings scheme is presented by the Hamiltonian
$a^{{\dagger}}\sigma+a\sigma^{{\dagger}}+a^{{\dagger}}a+\sigma^{{\dagger}}\sigma,$ (2)
@@
The weakness of the hydrogen bond causes its strong dependence on external conditions, in particular, on the ambient temperature. The temperature itself can be conventionally represented as a graph of the average number of phonons versus their frequency. In [50], the temperature effect on atoms within the Jaynes–Cummings model was represented as terms of the Hamiltonian of the form
$g(a_{\omega}^{\dagger}+a_{\omega})\sigma_{\omega}^{\dagger}\sigma_{\omega},$ (3)
where $a_{\omega}$ is the annihilation operator of the phonon of the selected mode $\omega$, $\sigma_{\omega}$ is the relaxation operator of the atom. This approach is extremely computationally expensive, since it requires explicit inclusion of phonons in the basic states of the model, which immediately causes a huge increase in the required memory and does not allow scaling the model to multi-molecular structures to study new collective effects. In addition, the case of atomic excitations is very different from Eq. (3), because their energy is several orders of magnitude greater than the energy of a single phonon, so that a term of the form Eq. (3) corresponds only to dephasing, but not to the direct interaction of phonons with matter.
The decoherence is described as caused by the influence of the medium using the QME with a decoherence factor in the form of phonon annihilation. This approach is simpler and more efficient than an independent a priori introduction of decoherence with a Gaussian factor in [10] and also allows for simple scaling.
Interest is only given to the frequency that most strongly affects the hydrogen bond. This influence can be represented as a term of the Hamiltonian of the form
$H_{int}=g_{hb}(a_{phn}^{{\dagger}}\sigma_{hb}+a_{phn}\sigma_{hb}^{\dagger}),$ (4)
where $\sigma_{hb}$ is the operator of the hydrogen bond formation accompanied by phonon emission, and $a_{phn}$ is the operator of phonon absorption leading to bond decay. Direct inclusion of such a term in the Hamiltonian is also unacceptable for us, since explicit phonons again arise in the basic states. An approximation in the form of the average number of phonons $n_{av}$ at the resonance frequency sensitive to hydrogen bonding is applied; since the operators of phonon annihilation and creation are proportional to the square root of their number $\sqrt{n_{av}}$, $H_{int}$ is replaced with the operator
$H^{\prime}_{int}=g_{0}\sqrt{n_{av}}(\sigma_{hb}^{\dagger}+\sigma_{hb}).$ (5)
Then the basic state will not contain an explicit number of phonons, and the model is scalable to more complex molecular systems.
@@
${\cal G}\left(T\right)=c\sum\limits_{n=0}^{\infty}e^{\frac{-\hbar\omega n}{KT}}|n\rangle\langle n|,$ (6)
where $\hbar$ is the reduced Planck constant, $\omega$ is the phononic mode, $n$ is the phonon number, and $T$ is the cavity temperature at a given frequency mode $\omega$, $K$ is the Boltzmann constant, and $c$ is the normalization factor. $\gamma_{in}$ and $\gamma_{out}$, which are the intensities of the inflow and dissipation of phonons into and out of a notional cavity around the molecules, are proposed to form a geometric progression with the denominator [51]
$\mu=\frac{\gamma^{in}}{\gamma^{out}}=e^{\frac{-\hbar\omega}{KT}},$ (7)
where $\gamma^{in}$ refers to the overall spontaneous inflow rate and $\gamma^{out}$ refers to the overall spontaneous emission rate. A stable temperature occurs only when $\gamma_{in}<\gamma_{out}$. Knowing these coefficients, or knowing the temperature $T$ directly, $n_{av}$ is obtained. In practice, the coefficients are found only by optimizing them using neural networks based on the experimental results. The temperature at a fixed photonic mode is determined in a similar way. Photonic modes relate to transformations of electron states, and phononic modes relate to proton oscillations.
@@ II.1 The target model
Interaction between two water molecules causes micro-oscillations of the hydrogen atom in one of the water molecules; it allows hydrogen bonds to form between water molecules, but this process does not break covalent bonds. The hydrogen bond formation mechanism is as shown in Fig. 1. In panel (a), when two $\rm{H}_{2}\rm{O}$ molecules that are moving freely are far apart, a hydrogen bond cannot form between the molecules. However, when the molecules move closer together, the hydrogen atom (proton) of one molecule and the oxygen atom of another molecule attract each other, and a stable hydrogen bond is formed. In panel (b), the hydrogen atom of one molecule, the “donor,” is attracted by the oxygen atom of another molecule, the “receptor,” to produce the tunneling effect. The proton tunneling in the normal state does not destroy the covalent bond between the proton and its parent molecule but only deforms it. In the case without considering electrons, the effect of some stretching of the covalent bond on the hydrogen bond formation is ignored, and it is considered a normal state with a covalent bond to avoid cluttering the target model. In the model with electrons, this difference is no longer ignored. Theoretically, in addition to 2 covalent bonds for the oxygen atom, there are actually 2 more hydrogen bonds with protons from neighboring molecules. The oxygen atom, and its covalent and hydrogen bonds form a tetrahedron.
@@
A hydrogen bond forms between water molecules when they get close enough, allowing one of the protons attached to one molecule to move between it and a nearby molecule while still keeping its bond. The qubit $|d\rangle_{dist}$ is introduced to describe the relative position of the proton in the system, and $d\in\{-1,0,1,2\}$. $dist=2$ indicates a significant distance between the molecules, rendering the formation of a hydrogen bond theoretically impossible. The state $dist=1$ means that the molecules are approaching each other at a critical distance, and although it does not allow the formation of a hydrogen bond, it allows the molecules to tunnel to the state $dist=0$, where the bond becomes possible. The state dist=1dist=-1 means the presence of a stable hydrogen bond; here, it is an irreversible process. However, researchers investigate the influence of temperature on the hydrogen bond formation, as a change in temperature triggers the inflow of phonons. Therefore, in this situation, the process becomes reversible, allowing for the breaking of hydrogen bonds.
Figure 1: (online color) Hydrogen bond formation mechanism. A scale model of two $\rm{H}_{2}\rm{O}$ molecules is shown in panel (a). The mechanism of ion attraction between hydrogen and oxygen atoms is shown in the form of a ball–and–stick model in panel (b). Brown circles (balls) represent oxygen atoms, red circles (balls) represent hydrogen atoms in an excited state, black solid lines stand for covalent bonds, and gray dotted lines represent ion attraction; especially, black dashed lines indicate stretched covalent bonds, and green balls in panel (a) represent protons in a ground state.
@@
$H_{i}=H_{i}^{basis}+H_{i}^{cogerence},$ (8)
where $H_{i}^{basis}$ is diagonal in the standard basis, and $H_{i}^{cogerence}$ is not diagonal in this basis. The term $H_{i}^{basis}$ is added to the total Hamiltonian in any case, and $H_{i}^{cogerence}$ is added only if the condition $A$ is satisfied. This remark about the conditional Hamiltonian will be valid for further models as well. This method achieves equality of rest energies for all terms in the general Hamiltonian. So the condition operator $Cond$ means only the inclusion of coherent terms in the Hamiltonian. For example, the proton (and, subsequently, the electron) jumps only if $d=0$, that is, when the molecules are close. Only at a large distance between the molecules is it as if it is “frozen” and does not lead to a change in the state of the proton (and, subsequently, the electron). And as soon as the molecules are close, the motions of these particles begin. Thus, the standard basis distinguishes itself from all other bases. In this work, we consider a single hydrogen bond between a pair of water molecules, which will be extended to more hydrogen bonds using Eq. (8) in the future.
Hydrogen bonding is a fundamental interaction in chemistry and biology, playing a crucial role in the structure and function of a wide range of molecules, from water and small organic compounds to large biomolecules like DNA and proteins. While there is no universally agreed-upon value for the hydrogen bond formation time, empirical studies suggest it typically occurs within the femtosecond to picosecond range, that is, from $10^{-15}$ to $10^{-12}$ seconds. This is due to the fact that the formation of hydrogen bonds involves the reorganization of electron clouds — a very fast quantum mechanical process. Time-scale measurements of the hydrogen bond dynamics can be performed using ultrafast spectroscopy techniques [52, 53, 54].